A plant has a pollution-control system shutdown once every 75 days, on average. What is the probability of at least one shutdown in the next year?

  1. 100.0%
  2. 1.33%
  3. 98.6%
  4. 99.3%
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

A. is correct.

x = 365 days

Lambda = 75 days

P(X>=1)

= 1- P(X=0)

= 1- (75^365)*(e^-75)/(365!)

= 1- 0

= 1 or 100%